OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x2 = −6x − 7?

hero (hero):

y = x^2-6x-7 y = x^2-7x+1x -7 y = x(x-7)+1(x-7) y = (x-7)(x+1)

hero (hero):

Set y = 0, then solve for x